steve invests $4070, part at 6% and part at 15% if the amount of interest earned at 6% was 4 times higher than the amount earned at 15%. how much did he invest at each rate.
Answer by Boreal(15235)

x=6% invested
4070-x=15% invested
.06x=4* 0.15(4070-x)
0.06x=4*(610.5-0.15x)
0.06x=2442-0.60x
0.66x=2442
x=3700
at 6%=$222
4070-x=$370
$370*0.15=$55.50
$55.50 *4=$222
$3700 at 6%
$370 at 15%